Given the following three functions: l(X) = x+4; h(x) = 1/x and g(X) = 2 - 11x. Derive the composite function h(l(g(X))).


1
Expert's answer
2022-01-31T18:07:31-0500

Solution:

"l(x) = x+4; h(x) = \dfrac 1x\ ;\ g(x) = 2 - 11x."

Now, "h(l(g(x)))=h(l(2-11x))"

"=h(2-11x+4)
\\=h(6-11x)
\\=\dfrac 1{6-11x}"